OpenStudy (anonymous):

For the graph shown, select the statement that best represents the given system of equations. 3y – x = 2 9y – 3x = 6 A.not enough information B.coincident C.consistent and independent D.inconsistent

jigglypuff314 (jigglypuff314):

If the system has at least one solution (one solution or infinitely many solutions), it is said to be consistent system. If a system has exactly one solution (a line intersects the other at some point), it is said to be consistent (has at least one solution) and independent (has exactly one solution). If the system has infinite number of solutions (all the points on the line will be the solution to that system), it is said to be consistent (has at least one solution) and dependent (has infinite number of solutions). If the system has no solution (parallel lines never intersect each other), it is said to be inconsistent system.

jigglypuff314 (jigglypuff314):

3y – x = 2 9y – 3x = 6 (divide through by 3) --> 3y - x = 2 therefore both equations are the same. therefore this has "infinitely many solutions"

OpenStudy (anonymous):

thx dude but the answer was C. consistent and independent, i got that one right anyways , thx for your help